# Disentangled and Self-Explainable Node Representation Learning

### v1

@@ Abstract
Node representations, or embeddings, are low-dimensional vectors that capture node properties, typically learned through unsupervised structural similarity objectives or supervised tasks. While recent efforts have focused on explaining graph model decisions, the interpretability of unsupervised node embeddings remains underexplored. To bridge this gap, we introduce DiSeNE (Disentangled and Self-Explainable Node Embedding), a framework that generates self-explainable embeddings in an unsupervised manner. Our method employs disentangled representation learning to produce dimension-wise interpretable embeddings, where each dimension is aligned with distinct topological structure of the graph. We formalize novel desiderata for disentangled and interpretable embeddings, which drive our new objective functions, optimizing simultaneously for both interpretability and disentanglement. Additionally, we propose several new metrics to evaluate representation quality and human interpretability. Extensive experiments across multiple benchmark datasets demonstrate the effectiveness of our approach.
@@ 1 Introduction
Self-supervised and unsupervised node representation learning [Ham20] provide a powerful toolkit for extracting meaningful insights from complex networks, making them essential in modern AI and machine learning applications related to network analysis [DLZ+24]. These methods offer flexible and efficient ways to analyze high-dimensional networks by transforming them into low-dimensional vector spaces. This transformation enables dimensionality reduction, automatic feature extraction, and the use of standard machine learning algorithms for tasks such as node classification, clustering, and link prediction [KSA19]. Furthermore, self-supervised node representations, or embeddings, enable visualization of complex networks and can be transferred across similar networks, enhancing understanding and predictive power in fields ranging from social networks to biological interactions.
Although widely adopted, unsupervised representation learning methods often face substantial challenges in terms of interpretability, necessitating complex and indirect approaches to understand what the learned embeddings actually represent [PKPA24, IKA20, GBH19]. This raises a critical question: What information do these embeddings encode?
While there has been a large body of work on explainable GNN models, limited attention has been given to embedding methods, which are the fundamental building blocks of graph based models. Existing efforts to explain embeddings are predominantly post-hoc [PKPA24, GBH19, KMA21, DJG18] and heavily dependent on the initial embedding techniques used. Some approaches [PKPA24] build upon existing literature for post-processing word embeddings [SPJ+18, CZ17], focusing on minimizing reconstruction errors of node embeddings using over-complete auto-encoders to improve sparsity. Other works [GBH19, KMA21, DJG18] focus solely on extracting meaningful explanations without addressing the learning process itself.
We propose DiSeNE (Disentangled and Self-Explainable Node Embedding), a framework that addresses these gaps by generating self-explainable unsupervised node embeddings. Our methodology is based on disentangled representation learning to encode node embeddings that are interpretable on a dimension-wise basis. Disentangled representation learning [WCT+22] aims to encode latent variables that represent separate factors of variation in data, ensuring that each latent dimension corresponds to a distinct, independent aspect.
@@
In graph data, node behaviour is strongly influenced by mesoscale structures like communities, which shape the network’s organization and drive dynamics [BBV08]. By leveraging disentangled representation learning, we capture these topological substructures more effectively, with each embedding dimension reflecting an independent graph unit (see Figure 3). We achieve this by introducing a new objective function to ensure structural disentanglement. Specifically, we optimize the embeddings so that each dimension is predictive of a unique substructure in the input graph. To avoid degenerate solutions, we incorporate an entropy-based regularizer which ensures that the mapped substructures are non-empty.
Our paradigm represents a shift in the language of explanations compared to the ones often considered when dealing with GNNs [YYGJ22]. Explainability for GNNs often involves understanding which parts of the local computation graph (nodes, edges) and node attributes significantly influence the model’s predictions [FKRA22, YBY+19, SEL+21]. On the other hand, the explanations that we aim to discover are inherently non-local, since they could involve mesoscale structures such as node clusters [PKPA24], usually not included in the GNN computational graph. To provide a comprehensive evaluation of the embeddings and uncover novel insights, we introduce several new metrics (for details refer to Section 3.3). For example, our overlap consistency measure (depicted in Figure 3) confirms that the physical overlap between the topological substructures identified as explanations aligns with the actual correlation of the corresponding embedding features or dimensions. This overlap offers insights into the interdependencies of node characteristics within the graph. Further, we attribute meaning to actual feature values by showing that the magnitude of node embedding entries corresponding to a dimension is correlated with the proximity (depicted in Figure 3) of the corresponding nodes to the “topological” subgraphs associated with that dimension. This measurement aids in understanding the closeness of nodes to different graph features, thereby enhancing spatial awareness within the graph structure.
To summarize we (i) formalize new and essential criteria for achieving disentangled and explainable node representations, offering a fresh perspective on explainability in unsupervised graph-based learning, (ii) introduce novel evaluation metrics to help quantifying the goodness of node representation learning in disentangled and explainable settings (iii) perform extensive experimental analyses to establish state-of-the-art results in self-explainable node feature learning. We release our code and data at https://github.com/simonepiaggesi/disene.
@@ 2 Preliminaries and Related Work
Given an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, node embeddings are obtained through an encoding function $\mathbf{h}:\mathcal{V}\rightarrow\mathbb{R}^{K}$ that map each node to a points of a $K-$dimensional vector space $\mathbb{R}^{K}$, where typically $D<<|\mathcal{V}|$. We denote the $K$-dimensional embedding of a node $v\in\mathcal{V}$ as $\mathbf{h}(v)=[h_{1}(v),\dots,h_{K}(v)]$, where $h_{d}(v)$ represents the value of the $d$-th feature of the embedding for node $v$. Alternatively, we can represent all node embeddings collectively as a matrix $\mathbf{H}(\mathcal{G})\in\mathbb{R}^{V\times K}$, where each entry $\mathbf{H}_{vd}=\mathrm{h}_{d}(v)$ corresponds to the $d$-th feature for node $v$. We can also refer to columns of such matrix, $\mathbf{H}_{:,d}$, as the dimensions of the embedding model space.
@@
Node embeddings are shallow encoding techniques, often based on matrix factorization or random walks [QDM+18]. Since the latent dimensions in these models are not aligned with high-level semantics [ŞUY+18, PDCM22], interpreting embeddings typically involves post-hoc explanations of their latent features [GBH19, KMA21]. Other works propose alternative methods to modify existing node embeddings, making them easier to explain with human-understandable graph features [PKPA24, SCER24]. From a different viewpoint, [SR24] explore how understandable are the embedded distances between nodes. Similarly, [DJG18] investigate whether specific topological features are predictable, and then encoded, in node representations.
@@
Graph Neural Networks (GNNs) [WPC+20] are deep models that operate via complex feature transformations and message passing. In recent years, GNNs have gained significant research attention, also in addressing the opaque decision-making process. Several approaches have been proposed to explain GNN decision process [YYGJ22], including perturbation approaches [YBY+19, YYW+21, FKRA22], surrogate model-based methods [VT20, HYT+22], and gradients-based methods [PKR+19, SLWL+20]. In parallel, alternative research directions focused on concept-based explanations, i.e. high-level units of information that further facilitate human understandability [MKSL21, XBG+23, MBK+22].
@@
Disentangled representation learning seeks to uncover and isolate the fundamental explanatory factors within data [WCT+22]. In recent years, these techniques have gained traction for graph-structured data [LWWX20, LWZ+21, YFSW20, FG24]. For instance, FactorGCN [YFSW20] disentangles an input graph into multiple factorized graphs, resulting in distinct disentangled feature spaces that are aggregated afterwards. IPGDN [LWWX20] proposes a disentanglement using a neighborhood routing mechanism, enforcing independence between the latent representations as a regularization term for GNN outputs. Meanwhile, DGCL [LWZ+21] focuses on learning disentangled graph-level representations through self-supervision, ensuring that the factorized components capture expressive information from distinct latent factors independently.
@@ 3.1 Core Objectives and Desiderata
In the context of self-supervised graph representation learning, we argue that learning self-explainable node embeddings amounts to reconstructing the input graph in a human-interpretable fashion. Traditionally, dot-product models based on NMF [YL13] and LPCA [CRR+24] decompose the set of graph nodes into clusters, where each entry of the node embedding vector represents the strength of the participation of the node to a cluster. In this scenario, the dot-product of node embeddings becomes intuitively understandable, as it reflects the extent of shared community memberships between nodes–thereby providing a clear interpretation of edge likelihoods. This concept is also related to distance encoding methods [LWWL20, KSN23], where a node feature $h_{d}(u)$ is expressed as a function of the node’s proximity $\zeta(u,\mathcal{S}_{d})=\mathrm{AGG}(\{\zeta(u,v),v\in\mathcal{S}_{d}\})$ to the anchor set $\mathcal{S}_{d}\subset\mathcal{V}$, using an aggregation function $\mathrm{AGG}$. Typically, distance encodings are constructed by randomly sampling anchor sets [YYL19], and used as augmented node features to enhance expressiveness and improve performance on downstream tasks.
@@
Inspired by this idea, our goal is to optimize self-supervised node embeddings encoded by a GNN function $\mathbf{h}:\mathcal{V}\rightarrow\mathbb{R}^{K}$ trained on the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, such that node features resemble non-random, structurally meaningful anchor sets, thus improving human-interpretability. To achieve this, we propose three key desiderata for learning general-purpose node representations: (i) structural faithfulness (ii), dimensional interpretability, and (iii) structural disentanglement. These desiderata serve as the foundational components of our approach, as detailed below.
Structural faithfulness.
@@
To implement the approach, we first train node embeddings in recovering the graph structure. We employ a random walk optimization framework based on the skip-gram model with negative sampling [HSS21]. The loss function for this framework is defined as:
$\mathcal{L}_{\text{rw}}=-\sum_{(u,v)\sim P_{rw}}\log\sigma\big{(}\mathbf{h}(u)^{\top}\mathbf{h}(v)\big{)}-\sum_{(u\textquoteright,v)\sim P_{n}}\log\sigma\big{(}-\mathbf{h}(u\textquoteright)^{\top}\mathbf{h}(v)\big{)},$
@@
where $\sigma(\cdot)$ is the sigmoid function, $P_{rw}$ is the distribution of node pairs co-occurring on random walks (positive samples), $P_{n}$ is a distribution over randomly sampled node pairs (negative samples), and $\mathbf{h}(u)^{\top}\mathbf{h}(v)$ represents the dot product between the embeddings of nodes $u$ and $v$. By optimizing this loss function, we encourage nodes that co-occur in random walks to have similar embeddings, effectively preserving the graph’s structural information in the embedding space.
@@
Given that our embeddings are structurally faithful–meaning they effectively encode the input graph’s structure–we should be able to interpret each embedding dimension in terms of the graph’s structural properties. We achieve this by attributing local subgraphs to different latent dimensions. Consider the likelihood of an edge $(u,v)$, defined as $\hat{y}(u,v;\mathbf{h})=\sigma\Big{(}\sum_{d=1}^{K}h_{d}(u)h_{d}(v)\Big{)}$. To understand how each dimension $d$ contributes to this likelihood, we compute the edge-wise dimension importance $\phi_{d}(u,v;\mathbf{h})$ as the deviation of the dimension-specific contribution from its average over all edges:
@@
Since the dot-product is a linear function $\sum_{d=1}^{K}\alpha_{d}h_{d}(u)h_{d}(v)+\beta$ with unitary coefficients $\alpha_{d}\equiv 1$ and zero intercept $\beta\equiv 0$, Eq. (1) corresponds to the formulation of LinearSHAP attribution scores [LL17], using the set of training edges as the background dataset. Essentially, the attribution function $\phi_{d}(u,v;\mathbf{h})$ indicates whether a specific dimension $d$ contributes positively to an edge’s likelihood. A positive attribution score means that the dimension increases the likelihood of predicting the edge. Leveraging this property, we generate dimension-wise explanations for the latent embedding model by collecting edge subsets with positive contributions:
@@
These self-explanations take the form of global edge masks $\mathbf{M}^{(d)}\in\mathbb{R}_{\geq 0}^{|\mathcal{V}|\times|\mathcal{V}|}$, where each entry is defined as $M^{(d)}_{uv}(\phi_{d};\mathbf{h})=\mathrm{max}\{0,\phi_{d}(u,v;\mathbf{h})\}$. By applying these masks to the adjacency matrix $\mathbf{A}$ through Hadamard product ($\odot$), we obtain $\mathbf{A}^{(d)}=\mathbf{A}\odot\mathbf{M}^{(d)}$. Each masked adjacency matrix $\mathbf{A}^{(d)}$ highlights the subgraph associated with dimension $d$. From these masked adjacency matrices, we construct edge-induced subgraphs $\mathcal{G}_{d}=(\mathcal{V}_{d},\mathcal{E}_{d})$, where $\mathcal{V}_{d}$ is the set of nodes involved in edges $\mathcal{E}_{d}$. These subgraphs act as anchor sets for the model, providing interpretable representations of how each embedding dimension relates to specific structural patterns within the graph. We will refer to edge-induced subgraphs computed as the aforementioned procedure as explanation subgraphs/substructures or topological components of the embedding model.
@@
To enhance the effectiveness of structurally faithful encodings, each dimension of the latent space should encode an independent structure of the input graph, effectively acting as an anchor subgraph. Inspired by community-affiliation models [YL13, YL12], we introduce a node affiliation matrix $\mathbf{F}\in\mathbb{R}^{{|\mathcal{V}|}\times K}$ that captures the association between each node $u\in\mathcal{V}$ and anchor subgraph $\mathcal{G}_{d}=(\mathcal{V}_{d},\mathcal{E}_{d})$. Specifically, each entry $\mathbf{F}_{ud}$ is proportional to the magnitude of predicted meaningful connections between node $u$ and other nodes in $\mathcal{G}_{d}$, expressed using the per-dimension attribution scores from Eq. (1):
$\mathbf{F}_{ud}(\{\mathbf{h}\})=\sum_{v\in\mathcal{V}_{d}}\phi_{d}(u,v;\mathbf{h})$ (3)
This aggregates the contributions of dimension $d$ to the likelihood of edges involving node $u$. To achieve structure-aware disentanglement, we enforce soft orthogonality among the columns of the affiliation matrix [^1] Note that $\mathbf{F}_{ud}(\{\mathbf{h}\})\equiv\sum_{v\in\mathcal{V}}h_{d}(u)h_{d}(v)-\frac{|\mathcal{V}|}{|\mathcal{E}|}\sum_{(u^{\prime},v^{\prime})\in\mathcal{E}}h_{d}(u^{\prime})h_{d}(v^{\prime})$ and assuming $|\mathcal{V}|<<|\mathcal{E}|$, the second term becomes negligible, allowing us to approximate $\mathbf{F}_{ud}(\{\mathbf{h}\})$ and reduce computational costs.. This ensures that different embedding dimensions capture independent structures, leading to nearly non-overlapping sets of predicted links for each dimension. We express the columns of the affiliation matrix as $\mathbf{F}_{:,d}$ and obtain the disentanglement loss function as:
@@
where $\cos\left(\mathbf{F}_{:,d},\mathbf{F}_{:,l}\right)$ denotes the cosine similarity between the $d$-th and the $l$-th columns of $\mathbf{F}$, and and $\delta_{d,l}$ is the Kronecker delta function ($1$ if $d=l$ , 00 otherwise). This approach enables us to obtain disentangled representations [WCT+22], where each dimension corresponds to a single underlying structural factor in the graph. Although it is possible to obtain disentanglement at more coarse-grained levels (e.g., with groups of dimensions), we focus on single-feature disentanglement because it inherently leads to dimension-wise interpretability.

### v2

@@ Abstract
Node embeddings are low-dimensional vectors that capture node properties, typically learned through unsupervised structural similarity objectives or supervised tasks. While recent efforts have focused on post-hoc explanations for graph models, intrinsic interpretability in unsupervised node embeddings remains largely underexplored. To bridge this gap, we introduce DiSeNE (Disentangled and Self-Explainable Node Embedding), a framework that learns self-explainable node representations in an unsupervised fashion. By leveraging disentangled representation learning, DiSeNE ensures that each embedding dimension corresponds to a distinct topological substructure of the graph, thus offering clear, dimension-wise interpretability. We introduce new objective functions grounded in principled desiderata, jointly optimizing for structural fidelity, disentanglement, and human interpretability. Additionally, we propose several new metrics to evaluate representation quality and human interpretability. Extensive experiments on multiple benchmark datasets demonstrate that DiSeNE not only preserves the underlying graph structure but also provides transparent, human-understandable explanations for each embedding dimension.
@@ 1 Introduction
Self-supervised and unsupervised node representation learning (Hamilton, 2020) provide a powerful toolkit for extracting meaningful insights from complex networks, making them essential in modern AI and machine learning applications for network analysis (Ding et al., 2024). These methods offer flexible and efficient ways to analyze high-dimensional networks by transforming them into low-dimensional vector spaces. This transformation enables dimensionality reduction, automatic feature extraction, and the deployment of standard machine learning algorithms for tasks such as node classification, clustering, and link prediction (Khosla et al., 2021). Moreover, unsupervised node representations, or embeddings, enable visualization of complex networks and can be transferred across similar networks, enhancing understanding and predictive power in fields ranging from social networks to biological systems. Despite their widespread utility, these approaches often face substantial challenges in terms of interpretability, typically relying on complex and post-hoc techniques to understand the latent information encoded within the embeddings (Piaggesi et al., 2024; Idahl et al., 2019; Gogoglou et al., 2019). This limitation raises a critical question: What information do these embeddings encode?
Despite a large body of research on explainable GNN models, embedding methods–the fundamental building blocks of graph-based systems–have received comparatively little attention. Most existing attempts to explain embeddings are predominantly post-hoc (Piaggesi et al., 2024; Gogoglou et al., 2019; Khoshraftar et al., 2021; Dalmia et al., 2018) and heavily dependent on the specific embedding techniques used. Some approaches (Piaggesi et al., 2024) extend methods from the post-processing of word embeddings (Subramanian et al., 2018; Chen & Zaki, 2017) by minimizing reconstruction errors through over-complete auto-encoders to improve sparsity. Other works (Gogoglou et al., 2019; Khoshraftar et al., 2021; Dalmia et al., 2018) focus solely on extracting meaningful explanations, without addressing the underlying embedding learning process.
We propose DiSeNE (Disentangled and Self-Explainable Node Embedding), a framework that addresses the interpretability gap in unsupervised node embeddings by generating inherently self-explainable node representations. In our approach, "self-explainable" means that each embedding dimension corresponds to a distinct subgraph that globally explains the structural information encoded within that dimension. These dimensional subgraphs highlight human-comprehensible functional components of the input graph, providing clear and meaningful insights. DiSeNE leverages disentangled representation learning, an approach that encodes latent variables corresponding to distinct factors of variation in the data (Wang et al., 2024), to produce node embeddings that are interpretable on a per-dimension basis.
@@
In graph data, node behaviour is strongly influenced by mesoscale structures such as communities, which shape the network’s organization and drive dynamics (Barrat et al., 2008). By leveraging disentangled representation learning, we capture these "topological" substructures more effectively, with each embedding dimension reflecting an independent graph unit (see Figure 3). We achieve this through a novel objective function that ensures structural disentanglement. Specifically, we optimize the embeddings so that each dimension is predictive of a unique substructure in the input graph. To avoid degenerate solutions, we incorporate an entropy-based regularizer that ensures the resulting substructures are non-empty and informative.
Our paradigm represents a shift in the language of explanations compared to the ones often considered when dealing with GNNs (Yuan et al., 2023). Explainability for GNNs often involves understanding which parts of the local computation graph (nodes, edges) and node attributes significantly influence the model’s predictions (Funke et al., 2023; Ying et al., 2019; Schnake et al., 2022). On the other hand, the explanations that we aim to discover are inherently non-local, since they could involve mesoscale structures such as node clusters (Piaggesi et al., 2024), usually not included in the GNN computational graph.
We provide a comprehensive evaluation of the embeddings and uncover novel insights by proposing several new metrics that capture the interplay between disentanglement and explainability (see Section 3.3 for details). For instance, our overlap consistency metric (illustrated in Figure 3) shows that the overlap of the topological substructures used as explanations matches the correlation among the corresponding embedding dimensions, providing insights into the interdependencies among node characteristics within the graph. Moreover, we attribute meaning to the embedding values by showing that the magnitude of a node’s entry in a specific dimension correlates with its proximity (as depicted in Figure 3) to the topological subgraphs associated with that dimension. This relationship enhances our understanding of the relative positioning of nodes with respect to different graph components, thereby enhancing spatial awareness of the graph structure.
Contributions are summarized as follows: (i) we formalize new and essential criteria for achieving disentangled and explainable node representations, offering a fresh perspective on interpretability in unsupervised graph-based learning; (ii) we introduce novel evaluation metrics to help quantifying the goodness of node representation learning in disentangled and explainable settings (iii) we perform extensive experimental analyses on synthetic and real-world data to establish state-of-the-art results in self-explainable node feature learning. We release our code and data at https://github.com/simonepiaggesi/disene.
@@ 2 Preliminaries and Related Work
Given an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, node embeddings are obtained through an encoding function $\mathbf{h}:\mathcal{V}\rightarrow\mathbb{R}^{K}$ that maps each node to a point in a $K$-dimensional vector space $\mathbb{R}^{K}$, where typically $D<<|\mathcal{V}|$. We denote the $K$-dimensional embedding of a node $v\in\mathcal{V}$ as $\mathbf{h}(v)=[h_{1}(v),\dots,h_{K}(v)]$, where $h_{d}(v)$ represents the value of the $d$-th feature of the embedding for node $v$. Alternatively, we can represent all node embeddings collectively as a matrix $\mathbf{H}(\mathcal{G})\in\mathbb{R}^{V\times K}$, where each entry $H_{vd}=h_{d}(v)$ corresponds to the $d$-th feature for node $v$. We can also refer to columns of such matrix, $\mathbf{H}_{:,d}$, as the dimensions of the embedding model space.
@@
Node embeddings are shallow encoding techniques, often based on matrix factorization or random walks (Qiu et al., 2018). Since the latent dimensions in these models are not aligned with high-level semantics (Senel et al., 2018; Prouteau et al., 2022), interpreting embeddings typically involves post-hoc explanations of their latent features (Gogoglou et al., 2019; Khoshraftar et al., 2021). Other works propose alternative methods to modify existing node embeddings, making them easier to explain with human-understandable graph features (Piaggesi et al., 2024; Shafi et al., 2024). From a different viewpoint, Shakespeare & Roth (2024) explore how understandable are the embedded distances between nodes. Similarly, Dalmia et al. (2018) investigate whether specific topological features are predictable, and then encoded, in node representations.
@@
Graph Neural Networks (GNNs) (Wu et al., 2021) are deep models that operate via complex feature transformations and message passing. In recent years, GNNs have gained significant research attention, also in addressing the opaque decision-making process. Several approaches have been proposed to explain GNN decision process (Yuan et al., 2023), including perturbation approaches (Ying et al., 2019; Yuan et al., 2021; Funke et al., 2023), surrogate model-based methods (Vu & Thai, 2020; Huang et al., 2023), and gradients-based methods (Pope et al., 2019; Sánchez-Lengeling et al., 2020). In parallel, alternative research directions focused on concept-based explanations, i.e. high-level units of information that further facilitate human understandability (Magister et al., 2021; Xuanyuan et al., 2023).
@@
Disentangled representation learning seeks to uncover and isolate the fundamental explanatory factors within data (Wang et al., 2024). In recent years, these techniques have gained traction for graph-structured data (Liu et al., 2020; Li et al., 2021; Yang et al., 2020; Fan & Gao, 2024). For instance, FactorGCN (Yang et al., 2020) disentangles an input graph into multiple factorized graphs, resulting in distinct disentangled feature spaces that are aggregated afterwards. IPGDN (Liu et al., 2020) proposes a disentanglement using a neighborhood routing mechanism, enforcing independence between the latent representations as a regularization term for GNN outputs. Meanwhile, DGCL (Li et al., 2021) focuses on learning disentangled graph-level representations through self-supervision, ensuring that the factorized components capture expressive information from distinct latent factors independently.
@@ 3.1 Core Objectives and Desiderata
In the context of unsupervised graph representation learning, we argue that learning self-explainable node embeddings amounts to reconstructing the input graph in a human-interpretable fashion. Traditionally, dot-product models based on NMF (Yang & Leskovec, 2013) and LPCA (Chanpuriya et al., 2023) decompose the set of graph nodes into clusters, where each entry of the node embedding vector represents the strength of the participation of the node to a cluster. In this scenario, the dot-product of node embeddings becomes intuitively understandable, as it reflects the extent of shared community memberships between nodes, thereby providing a clear interpretation of edge likelihoods. This concept is also related to distance encoding methods (Li et al., 2020; Klemmer et al., 2023), where a node feature $h_{d}(u)$ is expressed as a function of the node’s proximity $\zeta(u,\mathcal{S}_{d})=\mathrm{AGG}(\{\zeta(u,v),v\in\mathcal{S}_{d}\})$ to the anchor set $\mathcal{S}_{d}\subset\mathcal{V}$, using specific aggregation functions $\mathrm{AGG}$. Typically, distance encodings are constructed by randomly sampling anchor sets (You et al., 2019), and used as augmented node features to enhance expressiveness and improve performance on downstream tasks.
@@
Inspired by this idea, our goal is to optimize unsupervised node embeddings encoded by a GNN function $\mathbf{h}:\mathcal{V}\rightarrow\mathbb{R}^{K}$ trained on the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, such that node features resemble non-random, structurally meaningful anchor sets, thus improving human-interpretability. To achieve this, we propose three key desiderata for learning general-purpose node representations: (i) connectivity preservation, (ii) dimensional interpretability, and (iii) structural disentanglement. These desiderata serve as the foundational components of our approach, as detailed below.
Connectivity preservation.
@@
Ideally, node embeddings are constructed so that the geometric relationships in the low-dimensional space mirror the connectivity patterns of the original graph. Nodes with greater similarity in the network should be placed close to each other in the embedding space, and viceversa. To implement the approach, we train the node embedding function in recovering the graph structure. We employ a random walk optimization framework based on the skip-gram model with negative sampling (Huang et al., 2021). The loss function for this framework is defined as:
$\mathcal{L}_{\text{rw}}=-\sum_{(u,v)\sim P_{rw}}\log\sigma\big(\mathbf{h}(u)^{\top}\mathbf{h}(v)\big)-\sum_{(u’,v)\sim P_{n}}\log\sigma\big(-\mathbf{h}(u’)^{\top}\mathbf{h}(v)\big),$
@@
where $\sigma(\cdot)$ is the sigmoid function, $P_{rw}$ is the distribution of node pairs co-occurring on simulated random walks (positive samples), $P_{n}$ is a distribution over randomly sampled node pairs (negative samples), and $\mathbf{h}(u)^{\top}\mathbf{h}(v)$ represents the dot product between the embeddings of nodes $u$ and $v$. By optimizing this loss function, we encourage nodes that co-occur in random walks to have similar embeddings, effectively preserving the graph’s structural information in the embedding space.
@@
Embedding representations are multi-dimensional, with specific latent factors contributing in representing each dimension (e.g., social similarity, functional proximity, shared interests). A given edge between $u$ and $v$ may rely more heavily on certain dimensions. This perspective shows how local relationships (edges) are directly informed by the global structure of the embeddings, with each dimension contributing uniquely to reconstructing particular relationships. Given structure-preserving embeddings, meaning they effectively encode the input structure, we should be able to interpret each embedding dimension in terms of the graph’s topological structure. Specifically, our framework attributes "responsibility" for reconstructing a local relationship (an edge) to specific dimensions, based on their contribution to the edge likelihood or the probability of reconstructing that edge through the embeddings. We achieve this by assigning local subgraphs to different latent dimensions. Consider the likelihood of an edge $(u,v)$, defined as $\hat{y}(u,v;\mathbf{h})=\sigma\Big(\sum_{d=1}^{K}h_{d}(u)h_{d}(v)\Big)$. The likelihood score is obtained by applying the sigmoid to a linear combination of per-dimension products $h_{d}(u)h_{d}(v)$. Because the sigmoid is monotonic, each product’s sign and magnitude directly determine its influence on $\hat{y}$. Thus, we can interpret the edge likelihood by inspecting the raw products. This decomposition requires only linear additivity and it does not depend on any statistical relationships among the terms. To understand how each dimension $d$ contributes to this likelihood, we compute the edge-wise dimension importance $\phi_{d}(u,v;\mathbf{h})$ as the deviation of the dimension-specific contribution from its average over all edges:
@@
Since the dot-product is a linear function $\sum_{d=1}^{K}\alpha_{d}h_{d}(u)h_{d}(v)+\beta$ with unitary coefficients $\alpha_{d}\equiv 1$ and zero intercept $\beta\equiv 0$, Eq. (1) corresponds to the formulation of LinearSHAP attribution scores (Lundberg & Lee, 2017), using the set of training edges as the background dataset. Essentially, the attribution function $\phi_{d}(u,v;\mathbf{h})$ indicates whether a specific dimension $d$ contributes positively to an edge’s likelihood. A positive attribution score means that the dimension increases the likelihood of predicting the edge. If two dimensions are correlated, they may often push an edge with similar attributions, but each still provides an observable marginal deviation that we can measure. In other words, $\phi_{d}$ quantifies marginal responsibility, not exclusive responsibility; therefore it remains valid in the presence of correlations. Each dimension contributes to reconstructing specific edges or substructures, and together, these contributions create a unified representation of the entire graph. Thus, the global nature of embeddings emerges as a direct consequence of their contributions to local relationships. From this observation, we generate dimension-wise (global) explanations for the latent embedding model by collecting edge subsets with positive contributions:
@@
These self-explanations take the form of global edge masks $\mathbf{M}^{(d)}\in\mathbb{R}_{\geq 0}^{|\mathcal{V}|\times|\mathcal{V}|}$, where each entry is defined as $M^{(d)}_{uv}(\phi_{d};\mathbf{h})=\mathrm{max}\{0,\phi_{d}(u,v;\mathbf{h})\}$. By applying these masks to the adjacency matrix $\mathbf{A}$ through Hadamard product ($\odot$), we obtain $\mathbf{A}^{(d)}=\mathbf{A}\odot\mathbf{M}^{(d)}$. Each masked adjacency matrix $\mathbf{A}^{(d)}$ highlights the subgraph associated with dimension $d$. From these masked adjacency matrices, we construct edge-induced subgraphs $\mathcal{G}_{d}=(\mathcal{V}_{d},\mathcal{E}_{d})$, where $\mathcal{V}_{d}$ is the set of nodes involved in edges $\mathcal{E}_{d}$. These subgraphs act as anchor sets for the model, providing interpretable representations of how each embedding dimension relates to specific structural patterns within the graph. We will refer to edge-induced subgraphs computed as the aforementioned procedure (pseudo-code in Appendix E) as explanation subgraphs/substructures or topological components of the embedding model.
@@
To enhance the effectiveness of dimensionally interpretable encodings, each dimension of the latent space should encode an independent structure of the input graph, effectively acting as an anchor subgraph. Inspired by community-affiliation models (Yang & Leskovec, 2013; 2012), we introduce a node affiliation matrix $\mathbf{F}\in\mathbb{R}^{{|\mathcal{V}|}\times K}$ that captures the association between each node $u\in\mathcal{V}$ and anchor subgraph $\mathcal{G}_{d}=(\mathcal{V}_{d},\mathcal{E}_{d})$. Specifically, each entry $\mathbf{F}_{ud}$ is proportional to the magnitude of predicted meaningful connections between node $u$ and other nodes in $\mathcal{G}_{d}$, expressed using the per-dimension attribution scores from Eq. (1): $F_{ud}(\mathbf{h})=\sum_{v\in\mathcal{V}_{d}}\phi_{d}(u,v;\mathbf{h}).$ This aggregates the contributions of dimension $d$ to the likelihood of edges involving node $u$. To achieve structure-aware disentanglement, we enforce soft-orthogonality among the columns of the affiliation matrix [^1] Note that $\mathbf{F}_{ud}(\mathbf{h})\equiv\sum_{v\in\mathcal{V}}h_{d}(u)h_{d}(v)-\frac{|\mathcal{V}|}{|\mathcal{E}|}\sum_{(u^{\prime},v^{\prime})\in\mathcal{E}}h_{d}(u^{\prime})h_{d}(v^{\prime})$ and assuming $|\mathcal{V}|<<|\mathcal{E}|$, the second term becomes negligible, allowing us to approximate $\mathbf{F}_{ud}(\mathbf{h})$ and reduce computational costs.. This ensures that different embedding dimensions capture independent structures, leading to nearly non-overlapping sets of predicted links for each dimension. We express the columns of the affiliation matrix as $\mathbf{F}_{:,d}$ and obtain the disentanglement loss function as:
@@
where $\cos\left(\mathbf{F}_{:,d},\mathbf{F}_{:,l}\right)$ denotes the cosine similarity between the $d$-th and the $l$-th columns of $\mathbf{F}$, and and $\delta_{d,l}$ is the Kronecker delta function ($1$ if $d=l$ , 0 otherwise). This objective penalizes correlation between the edge importance attribution scores of different dimensions. By reducing redundancy, we ensure that each dimension contributes uniquely to the reconstruction of a small set of edges, promoting interpretability. This approach enables us to obtain disentangled representations (Wang et al., 2024), where embedding dimensions correspond to orthogonal latent factors of the input graph. Although higher-order disentanglement is possible (e.g., with groups of dimensions), we focus on single-feature disentanglement for achieving dimension-wise interpretability.